ys 2325 Exa ame: The farious Dr. x lives on a dark and lonely planet. He attempts to launch a 2000 k ne satellite order to take over his world. The kg. If radius wishes to launch the of the dark and x mean ontrolis 10e m and its mass is M 7.5 x satellite into a circular orbit 1500 km above the surface of the planet, (10 pts.) What would the period of the satellite be once it is in orbit 1500 km ab the surface? b) (10 pts.) How much energy will it take to launch the satellite from the ground im orbit? Neglect any non-gravitational effects.
Answer #1

m = 2000 kg, M = 7.5*10^23 kg,

R = 3.3*10^6 m , h = 1500 km

(a) From Keplers third law

T^2 = 4pi^2(R+h)^3/(GM)

T^2 = 4*3.14^2*(3300000+1500000)^3/(6.67*10^-11*7.5*10^23)

T = 9337.44 s

T = 9337.44/(60*60)

T = 2.6 hours

(b) potential energy of satellite PE = -GMm/2r

Required energy =W = change in potential energy

W = -GMm/2(R+h) + GMm/2R

W = (6.67*10^-11*7.5*10^23*2000/2)(-1/(3300000+1500000) +1/(3300000))

W = 4.74*10^9 J